undeadknight26 (undeadknight26):

I think is is A but not sure... The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (anonymous):

can someone hhelp me with my question?

undeadknight26 (undeadknight26):

@AwakeReality @hartnn @skullpatrol please help me...

hartnn (hartnn):

f(2) is -26 to get f(3), just plug in n = 3 in f(n) = -4•f(n -1) - 3•f(n - 2) what do u get ?

undeadknight26 (undeadknight26):

ok...

undeadknight26 (undeadknight26):

F(3) = -4 x 2 - 3 x 1 i get -11...

undeadknight26 (undeadknight26):

@hartnn???

hartnn (hartnn):

f(n-1)= f(3-1) = f(2) why did you put f(2) = 2 ??? f(2) is -26

undeadknight26 (undeadknight26):

well isn't is 3 - 1? so 4 x 2?

hartnn (hartnn):

its f(3-1) not 3-1

undeadknight26 (undeadknight26):

so -4 x 3 - 1 - 3 x 3 - 2?

undeadknight26 (undeadknight26):

i do not understand..

hartnn (hartnn):

why don't you use f(2) = -26 ?

undeadknight26 (undeadknight26):

where do i use that?

undeadknight26 (undeadknight26):

Im sorry i do not remember how to do this.

hartnn (hartnn):

f(n) = -4•f(n -1) - 3•f(n - 2) f(3) = -4 f(2) - 3 f(1) here f(2) = -26 f(1) =8

undeadknight26 (undeadknight26):

ohhhh!!! f(3) = -4 x -26 - 3 x 8 i got 128...

hartnn (hartnn):

no -4 x -26 - 3 x 8 = 80

undeadknight26 (undeadknight26):

oh yea i for got to subtract sorry...

undeadknight26 (undeadknight26):

so its the second one!!! lol im so stupid sorr for wasting so much of your time...

hartnn (hartnn):

yes, 2nd one. no problem :)